# Data assimilation for networks of coupled oscillators: Inferring unknown model parameters from partial observations

### v2

@@ 3.3.2 Complex network topologies
As an example, for an Erdős-Rényi graph $N=50$ nodes and coupling probability $p=0.1$, such that each pair of nodes are coupled with probability $p$, the mean degree is $\langle k\rangle=N(p-1)=4.5$. The equivalent ring topology has $r^{*}=2.45$. The two nearest integers are $r_{1}=2$ and $r_{2}=3$. From the method for ring topologies in the previous section we obtain $\lambda_{1}=0.627$ and $\lambda_{2}=0.460$. Substituting these values into (26), (27) yields $\lambda^{*}=0.539$, which is the value of $\lambda$ used in Section 4.1.3 for random Erdős-Renyi networks.
3.4 Comparison between matrix exponential localization and Gaspari-Cohn localization functions
@@
When considering localization on networks, there are many possible definitions of the “distance” between two nodes. Possibly the simplest definition is the shortest path length, that is, the distance $d(i,j)$ between nodes $i$ and $j$ is the length of the shortest path connecting $i$ to $j$. With this definition, one can, in principle, define the function $C$ from (28), and create a matrix $\tilde{L}_{ij}=C(d(i,j),c)$, where here $c$ is a characteristic length for meaningful correlations. However, this construction does not in general yield a correlation matrix, in particular, the matrix $\tilde{L}$ is not always positive semi-definite. Nevertheless, we can compare the matrix $\tilde{L}$ with $L$ obtained from (22). In Section 3.3 we noted that for ring topologies correlations appear to be spurious for node distances greater than $2$, so a choice $c\approx 2$ is natural. Fig. 7 shows the first rows of the respective localization matrices $L$ and $\tilde{L}$ for the ring topology with $N=50$ and $r=3$ using the matrix exponential construction (22) (black crosses) compared to the Gaspari-Cohn construction with shortest path length distance (red circles). There is very good agreement between the two constructions.
@@
for $n\ll N$ and with one parameters $\lambda$. However, one would have to check that $\sum_{i=0}^{n}\frac{\lambda^{i}}{i!}\eta^{i}$ is non-negative for all eigenvalues $\eta$ of $A$. Consideration of sparse localization matrices of this form is left as an avenue for future work.
@@ 4 Results
Each of our data assimilation trials are initialized with an initial ensemble ${\bm{X}}_{0}^{\rm a}$ at $t=0$ with $M=2N+1$ members (the minimum required for the covariance matrices to be full rank) and $\bm{\phi}_{0}^{\rm a}\sim{\mathcal{N}}(\bm{\phi}_{0}+\bm{\eta}_{\phi},\sigma_{\phi,1}^{2}I)$, such that $\bm{\phi}_{0}$ is the true value at $t=0$ and $\bm{\eta}_{\phi}\sim\mathcal{N}(0,\sigma_{\phi,2}^{2})\in\mathbb{R}^{N}$. In essence, $\sigma_{\phi,2}$ controls the perturbations of the ensemble means away from the truth, and $\sigma_{\phi,1}$ controls the spread of the ensemble around the ensemble means. The initial ensemble for the parameters $\bm{p}$ has the analogous form $\bm{p}_{0}^{\rm a}\sim{\mathcal{N}}(\bm{p}_{0}+\bm{\eta}_{p},\sigma_{p,1}^{2}I)$, such that $\bm{\eta}_{p}\sim\mathcal{N}(0,\sigma_{p,2}^{2})\in\mathbb{R}^{N}$. For data assimilation with the Kuramoto model we use $\sigma_{\phi,1}^{2}=\sigma_{\phi,2}^{2}=0.25$, and $\sigma_{\omega,1}^{2}=\sigma_{\omega,2}^{2}=0.25\,\sigma_{\omega}^{2}=0.025$, where $\sigma_{\omega}^{2}=0.1$ is the variance of the natural frequency distribution. For data assimilation with the theta neuron network we use $\sigma_{\phi,1}^{2}=\sigma_{\phi,2}^{2}=0.04$, and $\sigma_{\zeta,1}^{2}=\sigma_{\zeta,2}^{2}=0.04\,\sigma_{\zeta}^{2}=0.004$, where $\sigma_{\zeta}^{2}=0.1$ is the variance of the intrinsic firing parameter distribution.
@@ 4.1 Data assimilation for the Kuramoto model
In most of our analyses (all except in Section 4.1.2) we initialize the Kuramoto model with a uniformly random initial condition of phases $\phi_{i}(0)$ on $[0,2\pi)$. We then observe the transition of the system from an initially incoherent state to a synchronized state for sufficiently large coupling strength $\kappa$. Synchronization presents a challenge for all data-driven methods. In a fully synchronized state, phases co-rotate at a uniform frequency, which is effectively a stationary state in a co-rotating reference frame. Once synchronization occurs, the observations progressively decrease in value, since they yield little new information. In Section 4.1.2 we discuss DA applied to cases with an initially synchronized state. We show that DA is still able to make improvements on the estimates of the phases $\phi_{i}$ and natural frequencies $\omega_{i}$, but the gain in accuracy is less than that obtained from observing the transient which carries more dynamic information.

### v3

@@ 3.3.2 Complex network topologies
As an example, for an Erdős-Rényi graph $N=50$ nodes and coupling probability $p=0.1$, such that each pair of nodes are coupled with probability $p$, the mean degree is $\langle k\rangle=N(p-1)=4.5$. The equivalent ring topology has $r^{*}=2.45$. The two nearest integers are $r_{1}=2$ and $r_{2}=3$. From the method for ring topologies in the previous section we obtain $\lambda_{1}=0.627$ and $\lambda_{2}=0.460$. Substituting these values into (26), (27) yields $\lambda^{*}=0.539$, which is the value of $\lambda$ used in Section 4.14.1.3 for random Erdős-Renyi networks.
3.4 Alternative localization matrices
@@
When considering localization on networks, there are many possible definitions of the “distance” between two nodes. Possibly the simplest definition is the shortest path length, that is, the distance $d(i,j)$ between nodes $i$ and $j$ is the length of the shortest path connecting $i$ to $j$. With this definition, one can, in principle, define the function $C$ from (28), and create a matrix $\tilde{L}_{ij}=C(d(i,j),c)$, where here $c$ is a characteristic length for meaningful correlations. However, this construction does not in general yield a correlation matrix, in particular, the matrix $\tilde{L}$ is not always positive semi-definite, which may cause the EnKF to “blow up” due to the term $(HP^{\rm f}_{n}H^{T}+R)^{-1}$ in the Kalman gain (4). Nevertheless, we can compare the matrix $\tilde{L}$ with $L$ obtained from (22). In Section 3.3 we noted that for ring topologies correlations appear to be weak for node distances greater than $2$, so a choice $c\approx 2$ is natural. Fig. 7 shows the first rows of the respective localization matrices $L$ and $\tilde{L}$ for the ring topology with $N=50$ and $r=3$ using the matrix exponential construction (22) (black crosses) compared to the Gaspari-Cohn construction with shortest path length distance (red circles). There is very good agreement between the two constructions.
@@
for $n\ll N$ and with one parameter $\lambda$. However, one would have to check that $\sum_{i=0}^{n}\frac{\lambda^{i}}{i!}\eta^{i}$ is non-negative for all eigenvalues $\eta$ of $A$. Consideration of sparse localization matrices of this form is left as an avenue for future work.
@@ 4 Results
Each of our data assimilation trials are initialized with an initial ensemble ${\bm{X}}_{0}^{\rm a}$ at $t=0$ with $M=2N+1$ members and $\bm{\phi}_{0}^{\rm a}\sim{\mathcal{N}}(\bm{\phi}_{0}+\bm{\eta}_{\phi},\sigma_{\phi,1}^{2}I)$, such that $\bm{\phi}_{0}$ is the true value at $t=0$ and $\bm{\eta}_{\phi}\sim\mathcal{N}(0,\sigma_{\phi,2}^{2})\in\mathbb{R}^{N}$. In essence, $\sigma_{\phi,2}$ controls the perturbations of the ensemble means away from the truth, and $\sigma_{\phi,1}$ controls the spread of the ensemble around the ensemble means. The initial ensemble for the parameters $\bm{p}$ has the analogous form $\bm{p}_{0}^{\rm a}\sim{\mathcal{N}}(\bm{p}_{0}+\bm{\eta}_{p},\sigma_{p,1}^{2}I)$, such that $\bm{\eta}_{p}\sim\mathcal{N}(0,\sigma_{p,2}^{2})\in\mathbb{R}^{N}$. For data assimilation with the Kuramoto model we use $\sigma_{\phi,1}^{2}=\sigma_{\phi,2}^{2}=0.25$, and $\sigma_{\omega,1}^{2}=\sigma_{\omega,2}^{2}=0.25\,\sigma_{\omega}^{2}=0.025$, where $\sigma_{\omega}^{2}=0.1$ is the variance of the natural frequency distribution. For data assimilation with the theta neuron network we use $\sigma_{\phi,1}^{2}=\sigma_{\phi,2}^{2}=0.04$, and $\sigma_{\zeta,1}^{2}=\sigma_{\zeta,2}^{2}=0.04\,\sigma_{\zeta}^{2}=0.004$, where $\sigma_{\zeta}^{2}=0.1$ is the variance of the intrinsic firing parameter distribution. The ensemble size $M=2N+1$ is chosen since it is the smallest ensemble size ensuring that the covariance matrices have full rank. This eliminates any problems stemming from rank-deficiency of the covariances (see [52]) and allows us to focus exclusively on the effect of the topological information of the localization.
@@ 4.1 Data assimilation for the Kuramoto model
In most of our analyses (all except in Section 4.14.1.2) we initialize the Kuramoto model with a uniformly random initial condition of phases $\phi_{i}(0)$ on $[0,2\pi)$. We then observe the transition of the system from an initially incoherent state to a synchronized state for sufficiently large coupling strength $\kappa$. Synchronization presents a challenge for all data-driven methods. In a fully synchronized state, phases co-rotate at a uniform frequency, which is effectively a stationary state in a co-rotating reference frame. Once synchronization occurs, the observations progressively decrease in value, since they yield little new information. In Section 4.14.1.2 we discuss DA applied to cases with an initially synchronized state. We show that DA is still able to make improvements on the estimates of the phases $\phi_{i}$ and natural frequencies $\omega_{i}$, but the gain in accuracy is less than that obtained from observing the transient which carries more dynamic information.
@@ 5 Conclusion
We have shown the effectiveness of using the matrix exponential of the network adjacency matrix as a localization matrix for dynamics on networks, but future work should consider other types of localization matrices for networks. For instance, sparse localization matrices could be considered, which improve computational efficiency (see Section 3.4), or transform-then-localize schemes could be employed [24].
In recent years there has been a surge in data-driven methods to perform inverse problems, such as entropic outlier sparsification and entropy-optimal scalable probabilistic approximation [62, 63, 64]. It would be interesting to see if our localization method can be beneficial for these methods as well.
@@ Acknowledgements
Data availability
The code that implements our methodology, generates all data used, and processes the data to create the figures in the manuscript can be found at Zenodo: https://doi.org/10.5281/zenodo.15092531